OpenStudy (anonymous):

if A-B = 45 degree , show that (1+ tanA ) * (1 + tanB ) = 2tanA .

OpenStudy (vishweshshrimali5):

\[\large{A-B = \cfrac{\pi}{4}}\] \[\large{\implies \tan(A-B) = \tan \cfrac{\pi}{4} = 1}\] \[\large{\implies \cfrac{\tan A - \tan B}{1+\tan A \tan B} = 1}\] \[\large{\implies \tan A - \tan B = 1 + \tan A \tan B}\] \[\large{\implies \tan A = 1+\tan B + \tan A ~ \tan B}\tag{1}\]

OpenStudy (vishweshshrimali5):

Now: \[\large{(1+\tan A)(1+\tan B) = 1+\tan A + \tan B + \tan A ~\tan B}\] \[\large{= \tan A + (1+\tan B + \tan A ~ \tan B)}\tag{2}\]

OpenStudy (vishweshshrimali5):

Now from equation (1) use the value of : 1+ tan B + tan A * tan B = tan A

OpenStudy (vishweshshrimali5):

Put this value in equation (2)

OpenStudy (vishweshshrimali5):

you would get: tan A + (tan A) = 2 tan A

OpenStudy (vishweshshrimali5):

\[\large{\color{red}{\text{Hence Proved}}}\]

OpenStudy (vishweshshrimali5):

@miniash11 Any doubt ? :)

OpenStudy (anonymous):

umm...yes... how did u get the equation 2 ? by adding 1 on both sides ?

OpenStudy (vishweshshrimali5):

No no I did nothing like that I just collected `1 + tan B + tan A tan B` together inside `( )` brackets and left the `tan A` outside the brackets.

OpenStudy (vishweshshrimali5):

The only formula I used here is: \[\large{\boxed{\tan (x-y) = \cfrac{\tan x - \tan y}{1 + \tan x \tan y}}}\] in: \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 \[\large{A-B = \cfrac{\pi}{4}}\] \[\large{\implies \color{red}{\tan(A-B)} = \tan \cfrac{\pi}{4} = 1}\] \[\large{\implies \color{red}{\cfrac{\tan A - \tan B}{1+\tan A \tan B}} = 1}\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (vishweshshrimali5):

Any doubt ? :)

OpenStudy (anonymous):

yes sir, i did understand that but the only point im stuck at is (1+tanA)(1+tanB)=1+tanA+tanB+tanA tanB -> this line =tanA+(1+tanB+tanA tanB)(2)

OpenStudy (vishweshshrimali5):

Please don't call me sir. I am just a student :) Call me Vishwesh or Vishu

OpenStudy (vishweshshrimali5):

Okay lets see if I can help you out in that

OpenStudy (vishweshshrimali5):

first of all expand (1+tan A)(1 + tan B) What will you get ?

OpenStudy (anonymous):

okay vishu

OpenStudy (vishweshshrimali5):

:D Hi @dg2 How are you ?

OpenStudy (anonymous):

1 + tanB + tanA + 1 + tanAtanB ?

OpenStudy (vishweshshrimali5):

Why that 1 again ?

OpenStudy (vishweshshrimali5):